RC meeting and Run Plan, Saturday Oct. 7, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
- After the RC meeting yesterday MCC called to say beam would be down for ~1h.
- Beam was restored during swing shift and we performed the regular >30min beam down checkout, eventually reaching 45uA pol He3 production.
- There were beam charge asymmetry issues observed in Halls A and C. This was diagnosed and discussed in #daq on slack. The issue appears to have been fixed.
- During owl shift the communication to the oven temperature was lost and seemed to be stuck at 220. Bill was able to come in and fix the issue at 5am. A massive thank you to him for that!
- Owl shift continued relatively uninterrupted production running
- Day shift has so far continued this production running. We were also able to transition back to the GEnII-3stream daq config which includes the sbs gems this morning.
- There will be a maintenance day during Day shift on Wednesday 10/11
- If you want to get into the hall, please communicate with the RC to get this organized!
Data collected :
- SWING: 0.36 C
- OWL: 0.472 C
- DAY: 0. C (ongoing)
Hall A Subsystems Status
- Hodo HV readout bugged and displaying the VMon instead of VMon - VSet. Alarm Handler and HV GUI restarted many times today. HV values are fine though.
- BBCal 7.4 dead. Discussion on changing PMT base at next available time.
Hall access requests
If anyone wants to make an access, let the RC know!
Please check when your shifts are!
Please sign up for any open shifts!
Short Term Run plan
Hall B have asked when we next plan an IHWP change
- Tentative plan with Halls B and C to replace IHWP on sunday.
- Standard production running at 45uA, with NMRs every 3 hours.
List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)